Researchers at the Mayo Clinic have studied the effect of sound levels on patient healing and have found a significant association (louder hospital ambient sound level is associated with slower postsurgical healing). Based on the Mayo Clinic's experience, Ardmore Hospital installed a new vinyl flooring that is supposed to reduce the mean sound level (decibels) in the hospital corridors. The sound level is measured at five randomly selected times in the main corridor.


New Flooring Old Flooring
42 48
41 51
40 44
37 48
44 52

1

(a-1)

Choose the correct null and alternative hypotheses. Assume μ1 refers to the average sound level of the new flooring and μ2 refers to the average sound level of the old flooring.

H0: μ1μ2 = 0 vs. H1: μ1μ2 ≠ 0

H0: μ1μ2 ≤ 0 vs. H1: μ1μ2 > 0

H0: μ1μ2 ≥ 0 vs. H1: μ1μ2 < 0

References

2.

(a-2)

At α = .05, what is the decision rule under the assumption that the variances are equal?

Reject the null hypothesis if tcalc > –1.86 (8 df.)

Reject the null hypothesis if tcalc < –1.86 (8 df.)

3.

(a-3)

Calculate the test statistic. (Round your answer to 2 decimal places. A negative amount should be indicated with a minus sign.)

4.

(a-4) At α = .05, has the mean been reduced?


  (Click to select)RejectDo not reject H0, the mean (Click to select)has not beenhas been reduced.

5.

(b-1)

Now, we are going to test whether or not the two variances are equal, rather than merely assume that they are equal. So, we now need to specify a new set of hypotheses based on the variances themselves. Select the appropriate hypotheses.

H0: σ12 / σ22 = 1vs. H1: σ12 / σ22 ≠ 1.

H0: σ12 / σ22 ≠ 1vs. H1: σ12 / σ22 = 1.

6.

(b-2)

At α = .05, what is the decision rule?

Reject H0 if Fcalc < 9.60 or Fcalc > .1042. (d.f.1 = 4, d.f.2 = 4.)

Reject H0 if Fcalc > 9.60 or Fcalc < .1042. (d.f.1 = 4, d.f.2 = 4.)

7.

(b-3) What is the test statistic? Put the smaller of the two variances in the numerator when you do the calculation. The reason for doing this -- rather than putting the larger of the two variances in the numerator -- may become apparent when you read the explanation that accompanies the last question. (Round the test statistic value to 2 decimal places.)

8.

(b-4) At α = .05, has the variance changed?


  (Click to select)RejectDo not reject H0, the variance (Click to select)has has not  changed.
Answer #1

Solution:

Ardmore Hospital installed a new vinyl flooring that is supposed to reduce the mean sound level (decibels) in the hospital corridors. The sound level is measured at five randomly selected times in the main corridor.

Part 1)

Choose the correct null and alternative hypotheses. Assume μ1 refers to the average sound level of the new flooring and μ2 refers to the average sound level of the old flooring.

Since Ardmore Hospital installed a new vinyl flooring that is supposed to reduce the mean sound level (decibels) in the hospital corridors and μ1 refers to the average sound level of the new flooring and μ2 refers to the average sound level of the old flooring, thus this is a left tailed test.

Thus correct hypothesis are:

Option c) H0: μ1μ2 ≥ 0 vs. H1: μ1μ2 < 0

Part 2)

(a-2) At α = .05, what is the decision rule under the assumption that the variances are equal?

df = n1 + n2 - 2 = 5 + 5 - 2 = 8

t Table t 975 one-tail 0.50 0.25 0.20 0.15 0.10 0.05 0.025 two-tails1.00 0.50 0.400.30 0.20 0.10 0.05 9 cum. prob t95 1 0.000 1.000 1.376 1.963 3.078 6.114 12.71 2 0.000 0.816 1.061 386 1.886 2.20 4.303 3 0.000 0.765 0.978 1250 1638 2.53 3.182 40.000 0.741 0.941190 1533 2.132 2.776 50.000 0.727 0.920 156 1.476 2.015 2.571 6 0.000 0.718 0.906 1.134 1.440 1.43 2.447 71 0.000 0.711 0.896 1.119 1.415 1.895 2.365 08 1.397 1.8602.306

t = 1.860

Since this is left tailed test, t critical value is negative = -1.86

Thus correct option is:

Reject the null hypothesis if tcalc < –1.86 (8 df.)

Part 3)

(a-3)

Calculate the test statistic.

t = \frac{\bar{x_{1}}-\bar{x_{2}}}{\sqrt{S_{p}^{2}\times \left ( \frac{1}{n_{1}}+\frac{1}{n_{2}} \right )}}

where

S_{p}^{2}=\frac{(n_{1}-1)\times s_{1}^{2}+(n_{2}-1)\times s_{2}^{2}}{n_{1}+n_{2}-2}

s_{1}^{2}=\frac{\sum x_{1}^{2}-(\sum x_{1})^{2}/n_{1}}{n_{1}-1}

s_{2}^{2}=\frac{\sum x_{2}^{2}-(\sum x_{2})^{2}/n_{2}}{n_{2}-1}

Thus we need to make following table:

New Flooring x1 Old Flooring x2 x1^2 x2^2
42 48 1764 2304
41 51 1681 2601
40 44 1600 1936
37 48 1369 2304
44 52 1936 2704
\small \sum x_{1}=204 \small \sum x_{2}=243 \small \sum x_{1}^{2}=8350 \small \sum x_{2}^{2}=11849

Thus

s_{1}^{2}=\frac{\sum x_{1}^{2}-(\sum x_{1})^{2}/n_{1}}{n_{1}-1}

s_{1}^{2}=\frac{8350 -(204 )^{2}/5}{5-1}=\frac{8350-8323.2 }{4}=\frac{26.8 }{4}=6.7

s_{2}^{2}=\frac{\sum x_{2}^{2}-(\sum x_{2})^{2}/n_{2}}{n_{2}-1}

s_{2}^{2}=\frac{11849 -(243 )^{2}/5}{5-1}=\frac{11849-11809.8 }{4}=\frac{39.2 }{4}=9.8

Thus

S_{p}^{2}=\frac{(n_{1}-1)\times s_{1}^{2}+(n_{2}-1)\times s_{2}^{2}}{n_{1}+n_{2}-2}

S_{p}^{2}=\frac{(5-1)\times6.7+(5-1)\times 9.8}{5+5-2}

S_{p}^{2}=\frac{4 \times6.7+4 \times 9.8}{8}

S_{p}^{2}=\frac{26.8 +39.2 }{8}

S_{p}^{2}=\frac{66 }{8}

S_{p}^{2}=8.25

and

\bar{x}_{1}=\frac{\sum x_{1}}{n_{1}}=\frac{204}{5}=40.8

\bar{x}_{2}=\frac{\sum x_{2}}{n_{2}}=\frac{243}{5}=48.6

Thus t test statistic is:

t = \frac{\bar{x_{1}}-\bar{x_{2}}}{\sqrt{S_{p}^{2}\times \left ( \frac{1}{n_{1}}+\frac{1}{n_{2}} \right )}}

t = \frac{40.8 -48.6 }{\sqrt{8.25 \times \left ( \frac{1}{5}+\frac{1}{5} \right )}}

t = \frac{-7.8 }{\sqrt{8.25 \times \left ( \frac{2}{5}\right )}}

t = \frac{-7.8 }{\sqrt{3.3 }}

t = \frac{-7.8 }{1.8166 }

t = -4.29

Part 4) (a-4) At α = .05, has the mean been reduced?

Since tcal = -4.29 < t critical value = -1.86,

we reject H0. the mean has been reduced.